Hi! How do I solve this? Find the probability that no two people have the same birthday when the number of randomly selected people is 3? Thanks a lot!
Answer by venugopalramana(3286)

Hi! How do I solve this? Find the probability that no two people have the same birthday when the number of randomly selected people is 3? Thanks a lot!
THE FIRST PERSON CAN HAVE ANY ONE OF 365 DAYS AS HIS BIRTH DAY
THE II PERSON .......................364......................AS IT CANT BE SAME AS ABOVE
THE III ..............................363...................
SO NUMBER OF POSSIBILITIES TO MEET THE GIVEN REQUIREMENT = 365*364*363
TOTAL NUMBER OF POSSIBILITIES IS 365*365*365
PROBABILITY = 365*364*363/365*365*365 = 0.991796